OpenStudy (anonymous):

Factor x^2 + 64. Check your work.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (kc_kennylau):

not factorisable xD

OpenStudy (owlcoffee):

It is factorisable, but in the complex numbers.

OpenStudy (kc_kennylau):

(x+8i)(x-8i)

OpenStudy (skullpatrol):

Are you sure there is suppose to be a "+" not a "-" @mariacruz07 ?

OpenStudy (anonymous):

@skullpatrol yes thats exactly how the question looks.

OpenStudy (anonymous):

@kc_kennylau how did you get (x+8i)(x-8i)?

OpenStudy (kc_kennylau):

But if it's x^4+64 it can be factorized unto (x^2+x+8)(x^2-x+8)

OpenStudy (kc_kennylau):

x^2+64=x^2-(-64)=x^2-(8i)^2=(x+8i)(x-8i)

OpenStudy (anonymous):

@skullpatrol not really, thats why im confused. can you help me understand?

OpenStudy (kc_kennylau):

I believe that your textbook/website is wrong.

OpenStudy (skullpatrol):

kc_kennylau has given the answer, go to YouTube and look up the Khan Academy for complex numbers.

OpenStudy (anonymous):

@kc_kennylau wrong in what?

OpenStudy (anonymous):

@skullpatrol ok thank you, i will do that now.

OpenStudy (kc_kennylau):

But I can teach you complex numbers :D Complex numbers revolve around a concept, a number named \(i\), which is defined as: \[i=\sqrt{-1}\]

OpenStudy (anonymous):

@kc_kennylau thank you, you are so helpful!

OpenStudy (kc_kennylau):

The problem is do you wanna learn lol, coz at some extend it's difficult.

OpenStudy (anonymous):

@kc_kennylau yes i do wanna learn this!!

OpenStudy (kc_kennylau):

and we can deduce that \(\large i^2=-1\)

OpenStudy (kc_kennylau):

And complex numbers are numbers in the form of \(a+bi\), where \(a\) and \(b\) are real numbers.

OpenStudy (kc_kennylau):

So how would you find out the value of \((1+2i)+(3+4i)\)?

OpenStudy (anonymous):

would you substitute the i first?

OpenStudy (kc_kennylau):

No, i is not an unknown

OpenStudy (kc_kennylau):

well if substituting i as sqrt(-1) would make it easier, feel free to do so :)

OpenStudy (anonymous):

ok im not sure how to solve that.. sorry, im kinda lost.

OpenStudy (kc_kennylau):

oh, since you're new to this matter, let me give you an example first: \[(6+7i)+(5+6i)=6+7i+5+6i=11+13i\]

OpenStudy (kc_kennylau):

Terms with \(i\) can add together, terms without \(i\) can add together, but a term with \(i\) and a term without \(i\) cannot add together

OpenStudy (anonymous):

ok now im starting to get it

OpenStudy (kc_kennylau):

So can you do mine? :)

OpenStudy (anonymous):

yes give me a sec im gonna try it :)

OpenStudy (kc_kennylau):

ok :)

OpenStudy (anonymous):

4 + 6i ?

OpenStudy (kc_kennylau):

exactly, i can see that you're getting it now :D

OpenStudy (kc_kennylau):

What about \((3+4i)-(1+2i)\)? :)

OpenStudy (anonymous):

2 - 2i ? the "-" is getting me confused a bit

OpenStudy (kc_kennylau):

Just take out the brackets first :)

OpenStudy (anonymous):

oh so it would be 2 + 6i right?

OpenStudy (kc_kennylau):

ok, lemme give you an example: \[(7+6i)-(5+4i)=7+6i-5-4i=7-5+6i-4i=2+2i\]

OpenStudy (anonymous):

so im right, its 2 + 6i ?

OpenStudy (kc_kennylau):

nope :/ try to do it with steps :)

OpenStudy (kc_kennylau):

\[\begin{array}{ll}&(7+6i)-(5+4i)\\=&7+6i-5-4i\\=&7-5+6i-4i\\=&2+2i\end{array}\]

OpenStudy (kc_kennylau):

This is the example I gave you

OpenStudy (anonymous):

ok i did it like you did the example.. (3+4i) - (1 + 2i).... 3-1 + 4i - 2i = 2 + 2i ?

OpenStudy (kc_kennylau):

exactly :D

OpenStudy (kc_kennylau):

Now it's multiplication time... It may be a little bit difficult, so I'll start with this: \(2(3+4i)\)

OpenStudy (kc_kennylau):

Just distribute to get your answer :)

OpenStudy (anonymous):

6 + 8i?

OpenStudy (kc_kennylau):

yes :) Now do this \(2i(3+4i)\)

OpenStudy (kc_kennylau):

Again, just distribute :D

OpenStudy (kc_kennylau):

Don't forget that \(i^2=-1\) :)

OpenStudy (owlcoffee):

I'll take a little intervention here. I'll show her much slower the steps to take, in order to solve a substraction or sum of complex numbers. Now, you skipped one fundamental detail, the complex numbers behave as vectors, so you can't apply fully aritmethic properties because they won't work in many cases. your example was: \[(7+6i)-(5+4i)\] in order to turn the sustraction into a sum, I have to sum the opposite (just like vectors): \[(7+6i)+(-5-4i)\] so to solve the sum, you just englobe the real numbers and the complex part also: \[(7-5)+(6-4)i\] giving you a result of: \[2+2i\] Now, as a trick, always group the numbers without "i" and the numbers with "i" on a parenthesis, so it doesn't look confusing.

OpenStudy (anonymous):

6i + 8i^2 = 6i + 8(-1) = 6i - 8 ?

OpenStudy (kc_kennylau):

exactly :D You're doing very good here :)

OpenStudy (kc_kennylau):

Now I believe that you can calculate the value of (1+2i)(3+4i) :) Again, just distribute :D

OpenStudy (anonymous):

thank you, you are really teaching me in a way i understand. so i couldnt do this without your help!

OpenStudy (anonymous):

ok let me try this problem

OpenStudy (anonymous):

-5 + 10i or 10i - 5 ?

OpenStudy (kc_kennylau):

yes, and i'd prefer the first way of expression i.e. -5+10i :)

OpenStudy (kc_kennylau):

And I'm glad to see that you're getting it :)

OpenStudy (kc_kennylau):

Before we step into division, we must learn a thing called a "conjugate".

OpenStudy (kc_kennylau):

For any complex number \(a+bi\), its conjugate is \(a-bi\), and vice versa.

OpenStudy (anonymous):

so its basically the opposite sign?

OpenStudy (kc_kennylau):

yes :)

OpenStudy (kc_kennylau):

For example, the conjugate of 3+4i would be?

OpenStudy (anonymous):

3-4i

OpenStudy (kc_kennylau):

yep you are getting it :D

OpenStudy (anonymous):

:)

OpenStudy (kc_kennylau):

It's the first time I'm teaching someone something new, because my dream is to become a Math teacher :)

OpenStudy (kc_kennylau):

And I'm very glad to see that my teaching method actually works :D

OpenStudy (anonymous):

well you are very good at it, so im sure your dream will come true :)

OpenStudy (kc_kennylau):

Thanks :D

OpenStudy (kc_kennylau):

What's the conjugate of 5+2i?

OpenStudy (anonymous):

5-2i

OpenStudy (kc_kennylau):

Multiply 5+2i by its conjugate, what do you get? :)

OpenStudy (anonymous):

21?

OpenStudy (kc_kennylau):

yes :)

OpenStudy (kc_kennylau):

Noticed something strange?

OpenStudy (anonymous):

yes there is no complex number

OpenStudy (kc_kennylau):

exactly :D

OpenStudy (kc_kennylau):

Now we can step into division :)

OpenStudy (kc_kennylau):

Try to find the value of \(\dfrac{1+2i}{3+4i}\) :)

OpenStudy (kc_kennylau):

Hint: make the denominator real :)

OpenStudy (anonymous):

ok im stuck could u show me an example first?

OpenStudy (kc_kennylau):

\[\frac{5+6i}{7+8i}=\frac{(5+6i)(7-8i)}{(7+8i)(7-8i)}=\frac{83+2i}{113}=\frac{83}{113}+\frac2{113}i\]

OpenStudy (kc_kennylau):

And a little correction, (5+2i)(5-2i) is 29 not 21 :)

OpenStudy (kc_kennylau):

Do you get it? :/

OpenStudy (anonymous):

wait, how was it 29?

OpenStudy (kc_kennylau):

(5+2i)(5-2i) =25-10i+10i-4i^2 =25-4i^2 =25-4(-1) =25+4 =29

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

ok ima try the division problem now

OpenStudy (kc_kennylau):

Estoy correcto que hablas espanol? :)

OpenStudy (kc_kennylau):

De tu nombre lo veo yo

OpenStudy (anonymous):

si

OpenStudy (kc_kennylau):

...

OpenStudy (kc_kennylau):

por que no me dijiste tu?

OpenStudy (anonymous):

no sabia que hablabas español, i speak more english tho.

OpenStudy (kc_kennylau):

te gustaria si hablo ingles o espanol?